# Beyond-Diagonal RIS Architecture Design and Optimization under Physics-Consistent Models

### v1

@@ Abstract
Reconfigurable intelligent surface (RIS) is a promising technology for future wireless communication systems. Conventional RIS is constrained to a diagonal scattering matrix, which limits its flexibility. Recently, beyond-diagonal RIS (BD-RIS) has been proposed as a more general RIS architecture class that allows inter-element connections and shows great potential for performance improvement. Despite extensive progress on BD-RIS, most existing studies rely on simplified channel models that ignore practical electromagnetic (EM) effects such as mutual coupling and impedance mismatching. To address this gap, this paper investigates the architecture design and optimization of BD-RIS under the general physics-consistent model derived with multiport network theory in recent literature. Building on a compact reformulation of this model, we show that band-connected RIS achieves the same channel-shaping capability as fully-connected RIS, which extends existing results obtained for conventional channel models. We then develop optimization methods under the general physics-consistent model; specifically, we derive closed-form solutions for single-input single-output (SISO) systems, propose a globally optimal semidefinite relaxation (SDR)–based algorithm for single-stream multi-input multi-output (MIMO) systems, and design an efficient alternating direction method of multipliers (ADMM)–based algorithm for multiuser MIMO systems. Using the proposed algorithms, we conduct comprehensive simulations to evaluate the impact of various EM effects and approximations, including mutual coupling among RIS antennas and the commonly adopted unilateral approximation, on system performance.
@@ I Introduction
The concept of BD-RIS was first proposed in [4, 5], where the reconfigurable elements are interconnected through tunable impedances in groups (referred to as group-connected RIS) or all together (referred to as fully-connected RIS). In particular, the conventional RIS can be viewed as a special case of group-connected RIS in which each element forms an individual group. Thanks to the additional design flexibility provided by inter-element connections, BD-RIS is able to achieve higher channel gain [4] and sum-rate [5] compared to conventional RIS. In addition, unlike conventional RIS which only reflects signals, BD-RIS can work under multiple modes [5], including transmissive, reflective, and hybrid modes, and can also be deployed in a multi-sector setup to greatly enhance signal coverage [6]. Since then, BD-RIS has attracted significant research attention. Extensive studies have been devoted to designing efficient optimization algorithms [7, 8, 9, 10, 11], proposing new BD-RIS architectures that balance complexity and performance [12, 13, 14], assessing the performance of different BD-RIS architectures under non-ideal hardware [15, 16, 17, 18, 19], and exploring new application scenarios aided by BD-RIS [20, 21, 22]; see [23, 24] for more detailed summaries.
@@
  • Optimization under the general physics-consistent model. We study the optimization of RIS under the general physics-consistent model for different systems. For SISO systems, we show that the approach in [17] can be directly extended to obtain closed-form optimal solutions that maximize the channel gain of the general physics-consistent model. For MIMO systems, we propose a new semidefinite relaxation (SDR)–based algorithm that achieves the global optimal solution for the receive power maximization problem. To the best of our knowledge, the proposed algorithm is the first that guarantees global optimality, whereas no prior method achieves this even under the conventional channel model. For multiuser MIMO systems, we extend the algorithm in [10] for sum-rate maximization to the general physics-consistent model.
  • Performance assessment of RIS under the general physics-consistent model. Using our optimization framework, we study the impact of various EM effects and approximations through simulations. Our results reveal that stronger mutual coupling can enhance system performance in SISO, MIMO, and multiuser MIMO systems under Rayleigh fading channels, thereby extending the conclusion of [17] which considered only the SISO case. We also show that neglecting the mutual coupling among RIS elements in optimization leads to performance degradation. In contrast, the unilateral approximation remains highly accurate over a broad range, as long as the transmission distance is on the order of a few wavelengths. Therefore, this approximation can be reliably employed in practical system design.
@@ II-A The General Physics-Consistent RIS-Aided Channel Model
Consider the $N$-port network model for RIS-aided MIMO systems depicted in Fig. 1. Let $\mathbf{v}_{i}\in\mathbb{C}^{N_{i}}$, $i\in\{T,I,R\}$, denote the voltages at the transmit, RIS, and receive antenna array, respectively. The overall channel matrix $\mathbf{H}\in\mathbb{C}^{N_{R}\times N_{T}}$ is defined as the matrix that characterizes the linear relationship between the transmit and receive voltages, i.e.,
@@
where $\mathbf{Z}_{I}\in\mathbb{C}^{N_{I}\times N_{I}}$ is the impedance matrix of the reconfigurable impedance network. In this paper, we adopt the common assumption that the reconfigurable impedance network is reciprocal and lossless, which implies that $\mathbf{Z}_{I}$ is symmetric and purely imaginary [30, Chapter 4].
@@
Combining (1) – (5), it has been shown in [28, Section III] that
@@
Here, $\widetilde{\mathbf{Z}}=\left(\mathbf{I}+{\mathbf{Z}}_{0}\mathbf{Z}^{-1}\right)^{-1}$ and is partitioned as
@@ II-B Approximations and Approximate RIS-aided Channel Models
(A1) Unilateral approximation: The “unilateral approximation” assumes that the distances between the transmitter, RIS, and receiver are sufficiently large such that the electromagnetic effect from the receiving devices to the transmitting devices are negligible [29]. Under this assumption, we can set $\mathbf{Z}_{TI}=\mathbf{0},\mathbf{Z}_{TR}=\mathbf{0},\mathbf{Z}_{IR}=\mathbf{0}$ since these feedback links do not affect the channel matrix, which leads to the following approximate channel model [28, Eq. (64)]:
@@
(A2) Assumptions on the transmitter and receiver: The following two assumptions are commonly imposed at the transmitter and receiver for further simplification. First, we assume that the source impedances $\{Z_{T,n}\}_{1\leq n\leq N_{T}}$ at the transmitter and the load impedances $\{Z_{R,n}\}_{1\leq n\leq N_{R}}$ at the receiver are equal to the reference impedance $Z_{0}$, commonly set as $Z_{0}=50\;\Omega$, which gives $\mathbf{Z}_{R}=Z_{0}\mathbf{I}$ and $\mathbf{Z}_{T}=Z_{0}\mathbf{I}$. Second, we assume that the transmit and receive antennas are perfectly matched with no mutual coupling, and thus $\mathbf{Z}_{TT}=Z_{0}\mathbf{I}$ and $\mathbf{Z}_{RR}=Z_{0}\mathbf{I}.$ Under (A1) and (A2), the channel model simplifies to [28, Eq. (76)]
@@
(A3) Perfect matching and no mutual coupling among RIS: This assumption gives $\mathbf{Z}_{II}=Z_{0}\mathbf{I}$. Under (A1) – (A3), the channel model simplifies to [28, Eq. (77)]
@@
where $\mathbf{H}_{RT}:=\mathbf{S}_{RT}$, $\mathbf{H}_{RI}:=\mathbf{S}_{RI}$, and $\mathbf{H}_{IT}:=\mathbf{S}_{RT}$ denote the channels (scattering matrices) from transmitter to receiver, from RIS to receiver, and from transmitter to RIS, respectively.
@@ III-A An Explicit Expression for the General Physics-Consistent Model
In the above derivation, we have started from the $Z$-parameter-based channel model in (6) and arrived at an expression based on $Y$-parameter in (10), in which the effect of RIS is explicitly visible through the admittance matrix $\mathbf{Y}_{I}$. The $Y$-parameter turns out to be the most suitable representation for the general physics-consistent model, as it yields the simplest analytical expression. In addition, the $Y$-parameter enables flexible characterization of different RIS architecture; see Section IV-A and [12, 14] for details.
@@ III-B A Compact Form of the General Physics-Consistent Model
In this subsection, we derive a compact reformulation of (10). By the assumption that the RIS is lossless and reciprocal, $\mathbf{Y}_{I}$ is a purely imaginary symmetric matrix, which can be expressed as $\mathbf{Y}_{I}=\mathrm{i}\mathbf{B}_{I}$, where $\mathbf{B}_{I}\in\mathbb{R}^{N_{I}\times N_{I}}$ is the susceptance matrix satisfying $\mathbf{B}_{I}=\mathbf{B}_{I}^{T}$.
@@
Proposition 1 provides a significantly simplified representation of the model in (10). Notably, it mathematically shares the same form as (8), but the expression and physical meaning of each component are completely different in the two models. In (8), $\mathbf{H}_{RT}$, $\mathbf{H}_{RI}$, and $\mathbf{H}_{IT}$ are the corresponding submatrices of the scattering matrix $\mathbf{S}$ of the wireless channel, and $\bm{\Theta}$ is the scattering matrix of the RIS. In (12), $\bar{\mathbf{H}}_{RT}$, $\bar{\mathbf{H}}_{RI}$, and $\bar{\mathbf{H}}_{IT}$ are more intricate transformation of the admittance matrix $\mathbf{Y}$ and the load impedance $\mathbf{Y}_{R}$, and they do not exhibit an explicit relationship with the actual scattering matrix $\mathbf{S}$. The effect of RIS is fully captured by $\bar{\bm{\Theta}}$, which, due to the relation in (13), can be viewed as the scattering matrix of a “virtual” $N_{I}$-port network. The corresponding virtual susceptance matrix, $\bar{\mathbf{B}}_{I}$, is a linear transformation of the actual susceptance matrix of the RIS, as captured by (14). In fact, Proposition 1 implies that all existing channel models introduced in Section II can be represented in the form of (12), while the specific expressions of the components $\bar{\mathbf{H}}_{RT}$, $\bar{\mathbf{H}}_{RI}$, $\bar{\mathbf{H}}_{IT}$, and $\bar{\bm{\Theta}}$ vary across models.
@@ IV-B Optimal RIS Architecture under the General Physics-Consistent Model
Let $\mathcal{H}_{\text{fully}}$ and $\mathcal{H}_{\text{band},q}$ be the channel set achieved by fully-connected RIS and band-connected RIS with band-width $q$, respectively. Based on Proposition 1, $\mathcal{H}_{\text{fully}}$ and $\mathcal{H}_{\text{band},q}$ can be expressed as
@@
Proposition 2 indicates that band-connected RIS achieves the same channel shaping capability as fully-connected RIS in MIMO systems (with difference being a measure-zero set). Notably, the circuit complexity, measured by the number of required admittances, in the optimal band-connected RIS is $\mathcal{O}(N_{I}D)$, which is significantly lower than the $\mathcal{O}(N_{I}^{2})$ complexity of fully-connected RIS (given that $N_{I}\gg D$) [14]. As a special case of Proposition 2 with $N_{R}=1$, tree-connected RIS is optimal for MISO systems.

### v2

@@ Abstract
Reconfigurable intelligent surface (RIS) is a promising technology for future wireless communication systems. Conventional RIS is constrained to a diagonal scattering matrix, which limits its flexibility. Recently, beyond-diagonal RIS (BD-RIS) has been proposed as a more general RIS architecture class that allows inter-element connections and shows great potential for performance improvement. Despite extensive progress on BD-RIS, most existing studies rely on simplified channel models that ignore practical electromagnetic (EM) effects such as mutual coupling and impedance mismatching. To address this gap, this paper investigates the architecture design and optimization of BD-RIS under the general physics-consistent model derived with multiport network theory in recent literature. Building on a compact reformulation of this model, we show that band-connected RIS achieves the same channel-shaping capability as fully-connected RIS, which extends existing results obtained for conventional channel models. We then develop optimization methods under the general physics-consistent model; specifically, we derive closed-form solutions for single-input single-output (SISO) systems, propose a globally optimal semidefinite relaxation (SDR)–based algorithm for single-stream multi-input multi-output (MIMO) systems, and design an efficient alternating direction method of multipliers (ADMM)–based algorithm for multiuser MIMO systems. Using the proposed algorithms, we conduct comprehensive simulations to evaluate the impact of various EM effects and approximations. The results indicate that the commonly adopted unilateral approximation provides sufficient accuracy in RIS-aided systems and can therefore be readily adopted to simplify the channel model, whereas mutual coupling among RIS elements should be properly taken into account in channel modeling.
@@ I Introduction
The concept of BD-RIS was first proposed in [4, 5], where the reconfigurable elements are interconnected through tunable impedances in groups (referred to as group-connected RIS) or all together (referred to as fully-connected RIS). In particular, the conventional RIS can be viewed as a special case of group-connected RIS in which each element forms an individual group. Thanks to the additional design flexibility provided by inter-element connections, BD-RIS is able to achieve higher channel gain [4] and sum-rate [5] compared to conventional RIS. In addition, unlike conventional RIS which only reflects signals, BD-RIS can work under multiple modes [5], including transmissive, reflective, and hybrid modes, and can also be deployed in a multi-sector setup to greatly enhance signal coverage [6]. Since then, BD-RIS has attracted significant research attention. Extensive studies have been devoted to designing efficient optimization algorithms [7, 8, 9, 10, 11], proposing new BD-RIS architectures that balance complexity and performance [12, 13, 14], assessing the performance of different BD-RIS architectures under non-ideal hardware [15, 16, 17, 18, 19], and exploring new application scenarios aided by BD-RIS [20, 21, 22, 23]; see [24, 25] for more detailed summaries.
@@
  • Optimization under the general physics-consistent model. We study the optimization of RIS under the general physics-consistent model for different systems. For SISO systems, we show that the approach in [17] can be directly extended to obtain closed-form optimal solutions that maximize the channel gain of the general physics-consistent model. For single-stream MIMO systems, we carefully reformulate the receive power maximization problem and develop a novel semidefinite relaxation (SDR)-based algorithm. We prove that the proposed method attains the global optimum. To the best of our knowledge, this is the first approach that guarantees global optimality, whereas existing methods fail to do so even under the conventional channel model. For multiuser MIMO systems, we extend the algorithm in [10] for sum-rate maximization to the general physics-consistent model.
  • Performance assessment of RIS under the general physics-consistent model. Using our optimization framework, we study the impact of various EM effects and approximations through simulations. Our results reveal that stronger mutual coupling can enhance system performance in SISO, MIMO, and multiuser MIMO systems under Rayleigh fading channels, thereby extending the conclusion of [17] which considered only the SISO case. We also show that neglecting the mutual coupling among RIS elements in optimization leads to performance degradation. In contrast, the unilateral approximation remains highly accurate as long as the transmission distance is on the order of a few wavelengths. Hence, in typical RIS-aided wireless communication scenarios, this approximation can be reliably employed to simplify the channel model.
@@ II-A The General Physics-Consistent RIS-Aided Channel Model
Consider the $N$-port network model for RIS-aided MIMO systems depicted in Fig. 1. Let $\mathbf{v}_{i}\in\mathbb{C}^{N_{i}}$, $i\in\{T,I,R\}$, denote the voltages at the transmit, RIS, and receive antenna arrays, respectively. The overall channel matrix $\mathbf{H}\in\mathbb{C}^{N_{R}\times N_{T}}$ is defined as the matrix that characterizes the linear relationship between the transmit and receive voltages, i.e.,
@@
where $\mathbf{Z}_{I}\in\mathbb{C}^{N_{I}\times N_{I}}$ is the impedance matrix of the reconfigurable impedance network. In this paper, we adopt the common assumption that the reconfigurable impedance network is reciprocal and lossless, which implies that $\mathbf{Z}_{I}$ is symmetric and purely imaginary [31, Chapter 4].
@@
Combining (1) – (5), it has been shown in [29, Section III] that
@@
Here, $\widetilde{\mathbf{Z}}:=\left(\mathbf{I}+{\mathbf{Z}}_{0}\mathbf{Z}^{-1}\right)^{-1}$ and is partitioned as
@@ II-B Approximations and Approximate RIS-aided Channel Models
(A1) Unilateral approximation: The “unilateral approximation” assumes that the distances between the transmitter, RIS, and receiver are sufficiently large such that the electromagnetic effect from the receiving devices to the transmitting devices are negligible [30]. Under this assumption, we can set $\mathbf{Z}_{TI}=\mathbf{0},\mathbf{Z}_{TR}=\mathbf{0},\mathbf{Z}_{IR}=\mathbf{0}$ since these feedback links do not affect the channel matrix, which leads to the following approximate channel model [29, Eq. (64)]:
@@
(A2) Assumptions on the transmitter and receiver: The following two assumptions are commonly imposed at the transmitter and receiver for further simplification. First, we assume that the source impedances $\{Z_{T,n}\}_{1\leq n\leq N_{T}}$ at the transmitter and the load impedances $\{Z_{R,n}\}_{1\leq n\leq N_{R}}$ at the receiver are equal to the reference impedance $Z_{0}$, commonly set as $Z_{0}=50\;\Omega$, which gives $\mathbf{Z}_{R}=Z_{0}\mathbf{I}$ and $\mathbf{Z}_{T}=Z_{0}\mathbf{I}$. Second, we assume that the transmit and receive antennas are perfectly matched with no mutual coupling, and thus $\mathbf{Z}_{TT}=Z_{0}\mathbf{I}$ and $\mathbf{Z}_{RR}=Z_{0}\mathbf{I}.$ Under (A1) and (A2), the channel model simplifies to [29, Eq. (76)]
@@
(A3) Perfect matching and no mutual coupling among RIS: This assumption gives $\mathbf{Z}_{II}=Z_{0}\mathbf{I}$. Under (A1) – (A3), the channel model simplifies to [29, Eq. (77)]
@@
where $\mathbf{H}_{RT}:=\mathbf{S}_{RT}$, $\mathbf{H}_{RI}:=\mathbf{S}_{RI}$, and $\mathbf{H}_{IT}:=\mathbf{S}_{IT}$ denote the channels (scattering matrices) from transmitter to receiver, from RIS to receiver, and from transmitter to RIS, respectively.
@@ III-A An Explicit Expression for the General Physics-Consistent Model
Remark 1.
Under the unilateral approximation introduced in Section II-B, the model in (10) reduces to the approximate model $\mathbf{H}_{\text{app},1}.$ Specifically, with
$\mathbf{Z}_{\text{app}}=\left[\begin{matrix}\mathbf{Z}_{TT}&\mathbf{0}&\mathbf{0}\\ \mathbf{Z}_{IT}&\mathbf{Z}_{II}&\mathbf{0}\\ \mathbf{Z}_{RT}&\mathbf{Z}_{RI}&\mathbf{Z}_{RR}\end{matrix}\right],$
the corresponding approximate admittance matrix is given by
$\mathbf{Y}_{\text{app}}=\mathbf{Z}_{\text{app}}^{-1}=\left[\begin{matrix}\mathbf{Y}^{\text{app}}_{TT}&\mathbf{0}&\mathbf{0}\\ \mathbf{Y}_{IT}^{\text{app}}&\mathbf{Y}_{II}^{\text{app}}&\mathbf{0}\\ \mathbf{Y}_{RT}^{\text{app}}&\mathbf{Y}_{RI}^{\text{app}}&\mathbf{Y}_{RR}^{\text{app}}\end{matrix}\right],$
where $\mathbf{Y}_{ii}^{\text{app}}=\mathbf{Z}_{ii}^{-1},~i\in\{T,I,R\},$
$\mathbf{Y}_{IT}^{\text{app}}=-\mathbf{Z}_{II}^{-1}\mathbf{Z}_{IT}\mathbf{Z}_{TT}^{-1},~\mathbf{Y}_{RI}^{\text{app}}=\mathbf{Z}_{RR}^{-1}\mathbf{Z}_{RI}\mathbf{Z}_{II}^{-1},$
$\mathbf{Y}_{RT}^{\text{app}}=\mathbf{Z}_{RR}^{-1}(-\mathbf{Z}_{RT}+\mathbf{Z}_{RI}\mathbf{Z}_{II}^{-1}\mathbf{Z}_{IT})\mathbf{Z}_{TT}^{-1}.$
Here, we use the notation “app” to emphasize that the admittance matrix and each of its blocks are approximate, namely obtained by inverting the approximate impedance matrix $\mathbf{Z}_{\text{app}}$ with the unilateral approximation applied. Substituting $\mathbf{Y}_{\text{app}}$ into the general model in (10), we obtain
$\mathbf{H}_{\text{app}}\hskip-1.42271pt=\hskip-2.84544pt\left(\mathbf{Y}_{R}\hskip-0.85355pt+\hskip-0.85355pt\mathbf{Y}_{RR}^{\text{app}}\right)^{-1}\hskip-2.84544pt\left(-\mathbf{Y}_{RT}^{\text{app}}\hskip-0.85355pt+\hskip-0.85355pt\mathbf{Y}_{RI}^{\text{app}}\left(\mathbf{Y}_{I}\hskip-0.85355pt+\hskip-0.85355pt\mathbf{Y}_{II}^{\text{app}}\right)^{-1}\mathbf{Y}_{IT}^{\text{app}}\right).$
Further substituting the expressions of $\mathbf{Y}_{\text{app}}$ into the above equation yields $\mathbf{H}_{\text{app,1}}$.
In the above derivation, we have started from the $Z$-parameter-based channel model in (6) and arrived at an expression based on $Y$-parameter in (10), in which the effect of RIS is explicitly visible through the admittance matrix $\mathbf{Y}_{I}$. The $Y$-parameter turns out to be the most suitable representation for the general physics-consistent model, as it yields the simplest analytical expression. Since no constraints are imposed on $\mathbf{Y}_{I}$ in the above derivation, the channel model given in (10) does not rely on any specific characteristics of the reconfigurable impedance network. Hence, it also holds when the RIS is lossy or non-reciprocal. In the following subsection, we further derive a compact reformulation of (10) under the assumption that the RIS is lossless and reciprocal.
@@ III-B A Compact Form of the General Physics-Consistent Model
By the assumption that the RIS is lossless and reciprocal, $\mathbf{Y}_{I}$ is a purely imaginary symmetric matrix, which can be expressed as $\mathbf{Y}_{I}=\mathrm{i}\mathbf{B}_{I}$, where $\mathbf{B}_{I}\in\mathbb{R}^{N_{I}\times N_{I}}$ is the susceptance matrix satisfying $\mathbf{B}_{I}=\mathbf{B}_{I}^{T}$.
@@
Proposition 1 provides a significantly simplified representation of the model in (10). Notably, a channel model with the same mathematical structure was previously derived in [32] to analyze the effect of a decoupling network at the RIS with mutual coupling under the unilateral approximation. The same mathematical structure has also appeared in [17], where a tight upper bound on the channel gain achievable by BD-RIS has been derived under the unilateral approximation. In contrast to [32, 17], the model in Proposition 1 does not rely on the unilateral approximation and is therefore more general, and does not assume the presence of a decoupling network.
The model in Proposition 1 shares the same form as (8), but the expression and physical meaning of each component are completely different in the two models. In (8), $\mathbf{H}_{RT}$, $\mathbf{H}_{RI}$, and $\mathbf{H}_{IT}$ are the corresponding submatrices of the scattering matrix $\mathbf{S}$ of the wireless channel, and $\bm{\Theta}$ is the scattering matrix of the RIS. In (12), $\bar{\mathbf{H}}_{RT}$, $\bar{\mathbf{H}}_{RI}$, and $\bar{\mathbf{H}}_{IT}$ are more intricate transformation of the admittance matrix $\mathbf{Y}$ and the load impedance $\mathbf{Y}_{R}$, and they do not exhibit an explicit relationship with the actual scattering matrix $\mathbf{S}$. The effect of RIS is fully captured by $\bar{\bm{\Theta}}$, which, due to the relation in (13), can be viewed as the scattering matrix of a “virtual” $N_{I}$-port network. The corresponding virtual susceptance matrix, $\bar{\mathbf{B}}_{I}$, is a linear transformation of the actual susceptance matrix of the RIS, as captured by (14). In fact, Proposition 1 implies that all existing channel models introduced in Section II can be represented in the form of (12), thereby extending the results in [32, 17], while the specific expressions of the components $\bar{\mathbf{H}}_{RT}$, $\bar{\mathbf{H}}_{RI}$, $\bar{\mathbf{H}}_{IT}$, and $\bar{\bm{\Theta}}$ vary across models.
@@ IV-B Optimal RIS Architecture under the General Physics-Consistent Model
Let $\mathcal{H}_{\text{fully}}$ and $\mathcal{H}_{\text{band},q}$ be the channel sets achieved by fully-connected RIS and band-connected RIS with band-width $q$, respectively. Based on Proposition 1, $\mathcal{H}_{\text{fully}}$ and $\mathcal{H}_{\text{band},q}$ can be expressed as
@@
Proposition 2 shows that the difference between $\mathcal{H}_{\text{fully}}$ and $\mathcal{H}_{\text{band},q}$, i.e., the achievable channel sets of fully-connected and band-connected RIS, lies only in a low-dimensional subspace of $\mathcal{H}_{\text{fully}}$, which is negligible. Intuitively, this indicates that band-connected RIS achieves the same channel shaping capability as fully-connected RIS in MIMO systems, except for a very limited number of channel realizations, which will practically never occur. The circuit complexity, measured by the number of required admittances, in the optimal band-connected RIS is $\mathcal{O}(N_{I}D)$, which is significantly lower than the $\mathcal{O}(N_{I}^{2})$ complexity of fully-connected RIS (given that $N_{I}\gg D$) [14]. As a special case of Proposition 2 with $N_{R}=1$, tree-connected RIS is optimal for MISO systems.
@@
Figure 2: The curves of empirical culmultive distribution function (CDF) of relative approximation error defined in (19) for band-connected RIS with different band width $q$, where $N_{T}=N_{R}=4$.
[…]
$\mathbf{H}_{\text{band},q}^{(i)}=\arg\min_{\mathbf{H}\in\mathcal{H}_{\text{band,q}}}\|\mathbf{H}-\mathbf{H}_{\text{fully}}^{(i)}\|_{F}^{2}.$ (18)
We quantify the approximation accuracy using the empirical cumulative distribution function (CDF) of the relative error, defined as
$F_{q}(\epsilon)=\frac{1}{N_{\text{sample}}}\sum_{i=1}^{N_{\text{sample}}}\mathbf{1}\left({\|\mathbf{H}_{\text{band},q}^{(i)}-\mathbf{H}_{\text{fully}}^{(i)}\|_{F}}/{\|\mathbf{H}_{\text{fully}}^{(i)}\|_{F}}\leq\epsilon\right),$ (19)
where $\mathbf{1}(\cdot)$ denotes the indicator function, equal to $1$ if the condition holds and 0 otherwise. Fig. 2 depicts the curves of $F_{q}(\epsilon)$ for different $q$ values. As expected, $F_{q}(\cdot)$ increases with $q$. In particular, when $q$ takes the optimal value from Proposition 2, $F_{q}(\epsilon)=1$ for all $\epsilon\geq 0$. This demonstrates that, with probability one, any channel achievable by a fully-connected RIS can be exactly reproduced using a band-connected RIS (with $q$ given in Proposition 2), thereby validating Proposition 2.
Before concluding this section, we remark that if the MIMO channel transmits below its full DoF, the interconnections in the optimal architecture can further be reduced. Specifically, let $\hat{D}$ be the number of transmit data streams, the band width in the optimal band-connected RIS reduces to $q=2\min\{\hat{D},N_{I}/2\}-1$; see [14, Remark 5].